Find the answers to the following problems by providing line-by-line (step-by-step) solutions for each applicable question. You must show your work for questions 1-4 to get full credit for each question. Each question is worth 20 points. Please be sure to include the appropriate units as part of your final answer.

1. Mercury, the only metal that exists as a liquid at room temperature, melts at 234.3 K. Convert this melting point to degrees Celsius. 2. Mercury has a boiling point of 356.58°C. Convert this temperature to degrees Fahrenheit. 3. Mercury metal is poured into a graduated cylinder that holds exactly 22.5 mL. The mercury used to fill the cylinder weighs 306.0 g. From this information, calculate the density of mercury, in grams per cubic centimeter. 4. What is the density of mercury in pounds per cubic inch (lb/in3)? a. 1 lb = 454g, 1 in= 2.54 cm. 5. What are some of the physical and chemical properties of Mercury? Mention at least two of each.
1
Expert's answer
2018-04-27T14:17:08-0400
1. T(°C) = T(K) – 273.15
T(°C) = 234.3 – 273.15
T = -38.85 °C
2. T(°F) = T(°C)*9/5 + 32
T(°F) = 356.58*9/5 + 32
T = 673.84 °F
3. d = 306 g / 22.5 mL
d = 13.6 g/cm3
4. 13.6 g/cm3*2.543 cm3/in3 / 454 g/lb = 0.491 lb/in3
5. Physical properties of mercury include its density (13.6 g/cm3), color (silver), and melting point (-38.85 degrees Celcius).
Chemical properties include that it can be corrosive (will corrode aluminum), it forms amalgams with other metals, and is toxic (especially when found in organic compounds).